Solve for x:
2log5 10 - 1/4log5 x = log5 50
I know the answer is 16, but I don't know why or where to start on solving it. Thanks!

Answer by stanbon(57967)

Solve for x:
2log5 10 - 1/4log5 x = log5 50
log5 100 - log5 x^(1/4) = log5 50
log4[100/x^(1/5)] = log5 50
----------
100/x^(1/4) = 50
x^(1/4) = 100/50
x^(1/4) = 2
x = 2^4
x = 16
=============
Cheers,
Stan H.